OpenStudy (meehan98):

Probability Help, please! Harold flips a coin and rolls a standard number cube. Find the probability that the coin will show tails and the cube will show a seven. Write the probability as a fraction in simplest form.

OpenStudy (meehan98):

Doesn't a standard number cube have numbers 1-6?

OpenStudy (anonymous):

Yes it does

OpenStudy (meehan98):

So the probability of getting a 7 would be 0% So the answer is 0 ?

OpenStudy (meehan98):

@madhu.mukherjee.946 do you think you could help me with this, please?

OpenStudy (meehan98):

How is it 1/6 when there's not a number "7" on the cube, though?

OpenStudy (anonymous):

I never saw a number cube with the number '7' so I guess it's 0% probability?

OpenStudy (madhu.mukherjee.946):

yeah sorry

OpenStudy (madhu.mukherjee.946):

so total probablity will be 0

OpenStudy (madhu.mukherjee.946):

sorry @Meehan98 actually i am not well so i'm messing up with things

OpenStudy (meehan98):

No, it's fine! I just wanted to make sure I wasn't making a mistake.

OpenStudy (madhu.mukherjee.946):

okay

OpenStudy (meehan98):

I have a question on this one too because I feel like it would be 1/36 because P( rolling a six)=1/6 and P(spinning a purple)=1/6 and multiplied would equal 1/36, right?

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (anonymous):

Yes

OpenStudy (meehan98):

Okay, thank you both! Wish I could give each of you a medal!